What is the greatest amount of AlCl3 (in grams) that can be made with 114 grams of Al and 186 grams of Cl2? Which is the limiting reactant? Which reactant is in excess? (Al=26.98) (Cl=35.46) Al + Cl2 --> AlCl3

1
Expert's answer
2022-03-12T01:57:39-0500

The balanced equation is

2Al + 3Cl2 --> 2AlCl3


"n(Al)=\frac{114\ g}{26.98\ g/mol}=4.225\ mol"


"n(Cl_2)=\frac{186\ g}{70.92\ g/mol}=2.623\ mol"


According to the equation, 2 moles of Al react with 3 moles of Cl2. Actually, there are less moles Cl2 than Al, which obviously means that Cl2 is the limiting reactant.


"m(AlCl_3)=2.623\ mol(Cl_2)\times\frac{2\ mol(AlCl_3)}{3\ mol(Cl_2)}\times\frac{133.34\ g(AlCl_3)}{1\ mol(AlCl_3)}=233\ g"


Answer: Cl2 - limiting; Al - excess; 233 g AlCl3